OpenStudy (anonymous):

Differentiate function R(x)= (square root of 10)/ x^7

OpenStudy (anonymous):

i got square root of 10/ 7x^6 but it doesn't seem right

OpenStudy (anonymous):

use 10^1'2 * x^-7

OpenStudy (anonymous):

\[\frac{-7\sqrt{10}}{x^8}\]

OpenStudy (anonymous):

don't let the \[\sqrt{10}\] fool you. it is just a constant. like 3

OpenStudy (anonymous):

i dont undesrstand how you got -7 on numerator

OpenStudy (anonymous):

\[\frac{d}{dx}\frac{1}{x^7}=\frac{-7}{x^8}\]

OpenStudy (anonymous):

if you like \[\frac{1}{x^7}=x^{-7}\] so \[\frac{d}{dx}x^{-7}=-7x^{-8}=\frac{-7}{x^8}\]

OpenStudy (anonymous):

and the square root of ten just sits there. it is a number